OpenStudy (anonymous):

Segment DE is 5 cm long. Segment DF is 8 cm long. How long is segment EF? How do i solve it?

OpenStudy (triciaal):



OpenStudy (triciaal):

df = de +ef 8 = 5 + ef ef = 8-5 = 3

OpenStudy (gudden):

@triciaal .. bt it has not been given that if these points were that ways.... the distance EF can also be equals to 8+5 = 13... when D comes b/w E and F

OpenStudy (triciaal):

yeah i had considered that but the question said segment and is usually directional

OpenStudy (gudden):

. but sometimes...its not so...